OpenStudy (anonymous):

determine the transformation from the "parent function" then graph it. y=5-2^(x+3)

OpenStudy (precal):

5 is up five units - is reflection across x axis Is the parent funct 2^X?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

move \[y=2^x\] to the left 3 units. reflect it over the x - axis, then move it up 5 units

OpenStudy (precal):

Well if you use a graping calculator and graph the parent function, you can compare what the transformations are. ie y=2^x y=-2^x It will be obvious that the negative sign will reflect about the xaxis y=2^x has a doubling effect ie things will double while y=3^x will triple. I am not sure what that 3 is doing (My best guess is that it is moving the graph to the left three units.

OpenStudy (anonymous):

thanks for the help everyone, and im trying not to rely on the calculator so much cause i might not be able to use it in some classes for tests